Derive the equation a pentulam T.It depends on mass of the bob,length of the pentulam and acceleration due gravity?

1
Expert's answer
2021-12-08T19:48:41-0500

Period "T=" time to complete one oscillation

At Equilibrium "T_0=mg"

Restoring force on the pendulum mass displaced "F_e=-mgsin\theta"

For small displacements "sin\theta=\theta"

"\theta=\frac{displacement x}{Length l}=\frac{ x}{ l}"

Thus:

"Fe=-mg(\frac{x}{ l})"

From Newton's second law "a=\frac{F_e}{m}=- \frac{g x}{ l}"

Also for simple harmonic motion "a=-x\omega^2"

"-x\omega^2=-\frac{gx}{l}"

"\omega=\sqrt{g/l}"

Since period "T=2\pi/\omega"

"T=2\pi\sqrt{l/g}"